OpenStudy (anonymous):

Multiply. (x – 10)(x – 10) x2 – 100 x2 – 20x + 100 x2 – 20x – 100 x2 + 100

OpenStudy (anonymous):

x2 - 20x + 100

OpenStudy (anonymous):

x2 - 20x +100

OpenStudy (anonymous):

(x - 10) (x - 10) = x2 -10x - 10x + 100 = x2 - 20x + 100

OpenStudy (anonymous):

100+x^2-20x

OpenStudy (anonymous):

x^2 - 20x + 100

OpenStudy (anonymous):

@alanli123 is right. This is a formula: (x+y)^2 = x^2 + 2xy +y^2

OpenStudy (anonymous):

x * x = x^2 x * -10 = -10x x * -10 = -10x -10 * -10 = 100 Then add the -10x's and then you get x^2 - 20x + 100.

OpenStudy (anonymous):

Thank you guys! I chose him as the best answer because he replied first